OpenStudy (anonymous):

Using complete sentences, explain the limitations on the base of an exponential function

zepdrix (zepdrix):

The base? Hmmm. Well it can never be 0. Think about a^x if x=0 Then a^0 = 1 Hmmm there is no power that will give us a^x=0 (except in the limit). What else can we say, hmmmmm :o

OpenStudy (anonymous):

That's good! Thank you!